OpenStudy (anonymous):

what is the expression in factored form? 2x^2+16x+24

OpenStudy (anonymous):

@kimbllse777 :)

OpenStudy (idku):

\(\large\color{blue}{\displaystyle 2x^2+16x+24 }\) \(\large\color{blue}{\displaystyle \color{red}{2\cdot}x^2+\color{red}{2\cdot}8\times x+\color{red}{2\cdot}12 }\)

OpenStudy (emma.davenport):

(x+2)(x+6)

OpenStudy (idku):

\(\large\color{blue}{\displaystyle \color{red}{2}\left(x^2+8 x+12\right) }\)

OpenStudy (idku):

\(\large\color{blue}{\displaystyle \color{red}{2}\left(x^2+8 x+12\right) }\) \(\large\color{blue}{\displaystyle \color{red}{2}\left(x^2+2x+6x+12\right) }\) \(\large\color{blue}{\displaystyle \color{red}{2}\left(x(x+2)+6(x+2)\right) }\)

OpenStudy (idku):

and lastly, the final step.

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=2x%5E2%2B16x%2B24 does this help???

OpenStudy (anonymous):

no it did not @kimbllse777 but i think i got it thanks guys